OpenStudy (anonymous):

The volumes of two spheres are 729 in3 and 27 in3. What is the ratio of their radii, rounded to the nearest whole number @Directrix

Directrix (directrix):

If two solids are similar, the cube of the scale factor of the two solids is equal to the ratio of the volumes.

Directrix (directrix):

729/27 = (ratio of radii)^3 What is the cube root of 729/27?

OpenStudy (anonymous):

um i got 0.19 i know i probably did it wrong

Directrix (directrix):

(729/27) = (9*9*9)/3*3*3) Cube root is 9/3 = 3/1

Directrix (directrix):

Options?

OpenStudy (anonymous):

its 3:1 right?